20% of 600 is x less than 15% of 60x ?
Answer by Theo(13342)

your equation becomes:

20% of 600 = 15% of 60x - x
this is equivalent to:
.2*600 = .15*60x - x
simplify this to get:
120 = 9x - x
factor out the x on the right side of the equation to get:
120 = (9-1)*x
simplify to get:
120 = 8x
divide both sides of the equation by 8 to get:
120/8 = x
simplify to get:
15 = x

your solution is x = 15
plug that back into the original problem statement to see if the equation is true.
your original problem statement is:
20% of 600 is x less than 15% of 60x.
replace x with 15 and your problem statement becomes:
20% of 600 is 15 less than 15% of 60*15
simplify this to get:
120 is 15 less than 15% of 900
simplify this further to get:
120 is 15 less than 135 which makes 120 equal to 120 which is correct.

your solution is confirmed as good.
x = 15